40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5342     Entry time: Tue Sep 6 11:21:33 2011
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ITMX rehung (Friday) 

[Jenne, Katrin, Jamie]

I'm a bad kid, and forgot to elog my Friday morning work...

Bob gave me back ITMX after a 48hour bake at 80C + clean RGA scan Friday morning after coffee and doughnuts.  Katrin helped me put it back in the suspension wire. 

While I was leveling the optic (making sure the scribe lines on each side of the optic are at the same height off the table), Katrin cut some new viton for replacement EQ stops.  The optic was missing one lower earthquake stop (the one that Jamie noticed last week), and somehow one other rubber piece came out of the EQ stop on another lower screw while we were re-suspending the optic.  We put the new stops in, and then checked the balance of the test mass.

The oplev is still the HeNe laser that is leveled to the level optical table in the cleanroom.  The lever arm is ~1.5 meters, and over that distance the reflected beam was pointed "up" in pitch by ~1.5mm, which is less than one beam diameter of the HeNe.  This is well within our ability to correct using the OSEMs.

We then locked the test mass, and installed it in the chamber.  I approximately did the half-voltage centering of the OSEMs, leaving the fine-tuning to Kiwamu for after lunch. 

ELOG V3.1.3-